OpenStudy (anonymous):

Last quick question..please help! Solve 4 (^x + 7) = (^6x – 1). The options are: x ≈ 0.04 x ≈ 28.35 x ≈ 19.51 x ≈ 0.05

OpenStudy (anonymous):

I don't understand the other side. What is 6x-1 being raised to?

OpenStudy (anonymous):

OOPS sorry : 4(^x+7) = 6(^x-1).

OpenStudy (anonymous):

(x+7) * ln(4) = (x-1) * ln(6) x * ln(4) - x * ln(6) = -ln(6) - 7*ln(4) x (ln(4)-ln(6)) = (-ln(6) - 7*ln(4)) x = (-ln(6)-7*ln(4))/(ln(4)-ln(6))

OpenStudy (paxpolaris):

you should write 4^(x+7) = 6^(x-1) just to be clear ^=‘to the power of’

OpenStudy (anonymous):

ok,sry

OpenStudy (anonymous):

what answer did u get???

OpenStudy (anonymous):

Again, you got to do the calculation. I don't have a scientific calculator on my mac.

OpenStudy (anonymous):

i got 28.35!! :)

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

Sounds about right.